A certain ball always rebounds 2/3 of the distance from which it falls.If the ball is dropped from a height of 9 feet, and later it is observed rebounding to a height of 64/81 feet, then how many times did it bounce?


1
Expert's answer
2021-10-03T16:32:56-0400

"h_0=9"

"h_n=h_0(\frac{2}{3})^n=9(\frac{2}{3})^n"

"h_n=\frac{64}{81}\to 9(\frac{2}{3})^n=\frac{64}{81}\to (\frac{2}{3})^n=\frac{2^6}{3^6}"

Thus, n=6.

The ball bounced 6 times.